OpenStudy (anonymous):

Create a rational expression to be your game piece. You may choose from the list of factors below or make your own. There must be a variable term in both the numerator and denominator. (5x) (2x) (x + 4) (x – 5) (2x + 1) (3x + 5)

OpenStudy (anonymous):

@SolomonZelman

OpenStudy (solomonzelman):

I don't really get the question, am I making up any function?

OpenStudy (anonymous):

no, a rational expression

OpenStudy (anonymous):

like the one before

OpenStudy (solomonzelman):

Just any random rational expression?

OpenStudy (anonymous):

using two of the terms listed under

OpenStudy (solomonzelman):

And this would be just a rational expression? I mean it doesn't have to represent anything? just making up a math problem, right?

OpenStudy (anonymous):

yep

OpenStudy (solomonzelman):

\[\huge\color{red}{\frac{x^2-3x-4}{x^2+3x+2} \over \frac{x^2-2x-8}{x-1}}\]

OpenStudy (solomonzelman):

That's all?

OpenStudy (anonymous):

no lol you have to use the terms listed under the question,

OpenStudy (solomonzelman):

To make it? Damn, I used so other terms....

OpenStudy (anonymous):

LOL

OpenStudy (solomonzelman):

\[\huge\color{blue}{\frac{5x^2+20x}{2x^2+8x} \over \frac{x^2-x-20}{4x^2+2x}} \]

OpenStudy (solomonzelman):

Then what?

OpenStudy (anonymous):

choose from the factors listed in the question bro!

OpenStudy (solomonzelman):

I did, and made the rational expression from them in blue, above.

OpenStudy (anonymous):

ohhhh ok i see

OpenStudy (anonymous):

the next part is: Explain to the game master how to add your rational expression to the one on the correct space. Use complete sentences.

OpenStudy (solomonzelman):

add your rational expression to the one on the correct space?

OpenStudy (solomonzelman):

ID get that at all (sorry)

OpenStudy (anonymous):

oh my bad

OpenStudy (anonymous):

here maybe this will help

OpenStudy (anonymous):

First you will need to create your own rational expression. This will be your “game piece.” You will also need a coin with two different sides to keep track of “heads” versus “tails.”

OpenStudy (solomonzelman):

how does the part with coin relate to the rational expression?

OpenStudy (anonymous):

oh crap the lesson is corrupted. :( i can't give the info we need to do the problems

OpenStudy (anonymous):

ill have to contact my teacher. sorry !!

OpenStudy (solomonzelman):

It's fine, hope you get it.....

OpenStudy (anonymous):

ai

OpenStudy (anonymous):

heres a badge anyways :3

OpenStudy (anonymous):

are you good at physics?

OpenStudy (solomonzelman):

Not at all.

OpenStudy (solomonzelman):

I am worse at science than at anything.

OpenStudy (anonymous):

oh ok

OpenStudy (anonymous):

i just noticed your almost a champion!!

OpenStudy (solomonzelman):

Not really, I need lots of points to be champion, but ty!

OpenStudy (anonymous):

when i met you you were at 84 i believe?

OpenStudy (solomonzelman):

Something like that.